OpenStudy (anonymous):

A football field goal is kicked from 30 yards out (1 yard = 3 ft) at an angle of 32. The uprights of the goal post are 10 ft of the ground. What is the minimum speed required to just make a field goal?

OpenStudy (anonymous):

y=x tan\[\theta\] - .5*g*((vcos\[\theta\])^-2)*x^2 10=30tan32 -.5* 32.5 * (v cos32)^-2 * 30^2 .5* 32.5 * 30^2=.624*3 14625= (v cos32)^2 120.9=v*cos32 v=141.5 feet/s^2 may be...

OpenStudy (zion):

u cant kick a football field goal, its too heavy, it hurts. LMAO